OpenStudy (anonymous):

Find all values of x in the interval [0, 2π] that satisfy the equation. 3 sin(2x) = 3 cos(x)

OpenStudy (cwrw238):

sin 2x = 2 sinx cos x so: 3* 2 sinx cos x - 3 cos x = 0 3cos x ( 2 sin x - 1) = 0 cos x = 0 or 2sin x - 1 = 0 cos x = 0 or sin x = 1/2 for cos x = 0, one result is x = pi/2 for sin x = 1/2 one result is x = pi/6 can you find the others?

OpenStudy (anonymous):

@cwrw238 great job but it would appreciated if you have the step \[x= \cos^{-1}(0), and then say x = \pi/2, 3\pi/2,... \] as cos is 0 for pi/2 and 3pi/2 similarly \[x=\sin^{-1} (1/2)=\pi/6,(5\pi/6),...\] as SIN is positive in second quadrant

OpenStudy (anonymous):

HELLO you have cos(pi/2-2x)=sin(2x)......

OpenStudy (cwrw238):

????